les and gil drove a total of 733 miles in 11.6 hours. les drove the first part of the trip and averaged 55 mph. gil drove the remainder of the trip and averaged 65 mph. how long did les drive?
Answer by stanbon(48551)

les and gil drove a total of 733 miles in 11.6 hours.
les drove the first part of the trip and averaged 55 mph.
------------------------
Les DATA:
distance = x miles ; rate = 55 mph ; time = x/55 hrs.
-------------------------
gil drove the remainder of the trip and averaged 65 mph.
Gil DATA:
distance = 733 -x miles ; rate = 65 mph ; time = d/r = (733-x)/65 hrs
---------------------------
EQUATION:
time + time = time
x/55 + (733-x)/65 = 11.6
65x + 55(733-x) = 55*65*11.6
10x + 55*733 = 41470
10x = 1155
x = 115.5 miles (distance Les drove)
733-115.5 = 617.5 miles ( distance Gil drove)
================
Cheers,
Stan H.
how long did les drive?